23- 7(x+3) +5x=10
For my son's homework and i am stuck.

Found 2 solutions by checkley77, ly12603:
Answer by checkley77(12844)

23- 7(x+3) +5x=10
23-7x-21+5x=10
-2x=10-23+21
-2x=8
x=8/-2
x=-4 ans.
Proof:
23-7(-4+3)+5*-4=10
23-7*-1-20=10
23+7-20=10
30-20=10
10=10

Answer by ly12603(44)

Step 1: 23-7(x+3)+5x=10 <-- step one distribute
answer: 23-7x-21+5x=10
step 2: 2-7x+5x=10
step 3: 2-2x=10
step 4: -2x=8 <---divide by negative 2
step 5: x= -4